# Hanbury Brown-Twiss interference with massively parallel spectral multiplexing for broadband light

### v1

@@ Abstract
Two-photon interference in the spectral domain is a powerful resource for quantum technologies, enabling both precision measurements and scalable entanglement distribution. Here, we report the first demonstration of massively parallel, wavelength-resolved photon bunching, revealing Hanbury Brown–Twiss correlations across 100 independent spectral channels. These observations are enabled by a fast, data-driven single-photon spectrometer that achieves 40 pm spectral and 40 ps temporal resolution over a 10 nm bandwidth, allowing simultaneous access to spectro-temporal photon correlations without the need for narrowband filtering. This approach preserves photon flux while enabling high-dimensional quantum interference measurements across a broad spectrum. Our results establish frequency-multiplexed two-photon interference as a scalable and throughput-efficient platform for quantum-enhanced photonic technologies, with applications ranging from stellar intensity interferometry to broadband entanglement swapping in quantum networks.
@@ 1 Introduction
Two-photon correlation analysis stands as a cornerstone of contemporary quantum optics, tracing its origins back to foundational studies by Glauber [1, 2]. This powerful method has found diverse applications across numerous fields, including super-resolution microscopy [3, 4], quantum communication protocols [5], two-photon resonance fluorescence [6, 7], stellar intensity interferometry (SII) [8, 9, 10, 11, 12], and its advanced phase-sensitive variants [13, 14, 15, 16]. Central to these applications are distinct correlation signatures between photons arising from fundamental quantum interference effects, predominantly characterized by the indistinguishability of photons as in the Hanbury Brown-Twiss (HBT) [17, 18] and Hong-Ou-Mandel (HOM) [19] phenomena. In this study, we directly measure and quantify photon indistinguishability by employing a fast spectrometer combined with precise spectral and temporal data analysis conducted during post-processing as schematically presented in Figure 1. For the first time, we unambiguously observe the HBT effect of photon bunching in broadband light simultaneously across 100 spectral bins in a continuous 10 nm range, establishing a new methodology for massively multiplexed correlation sensing over an extensive frequency range, with the potential of substantially improving the resolution, sensitivity, and scalability for quantum optical measurements.
@@
Another application for the spectrometer with excellent spectral and temporal resolutions is resonance fluorescence, where precise two-photon interference and correlation measurements can yield groundbreaking insights into quantum optical processes in atomic and molecular physics. Spectral correlation patterns observed in resonance fluorescence typically manifest as distinct photon bunching lines and anti-bunching circles, unveiling quantum behaviors at off-resonant frequencies [25, 6].
@@
Previous attempts to measure the frequency-multiplexed HBT effect were constrained by either limited temporal or spectral resolution or by a narrow frequency range with few spectral bins. Earlier investigations of spectrally resolved HOM interference employed slower cameras, utilizing nanosecond-scale resolution to identify temporal coincidences in photon pairs. Temporal coherence was usually maintained by precisely delaying one photon relative to the other using mechanical stages, effectively enabling sub-picosecond precision and thereby providing good visibility for HOM interference [29, 30, 31]. More recent efforts, such as those described in [32, 33, 34], achieved significantly improved timing resolution, 20–40 ps. However, these studies still faced limitations, with only a small number of spectral channels, seven channels in [32] and five narrow neon spectral lines in [34], or restricted spectral coverage, less than 0.3 nm, when investigating two-photon transitions in polariton condensates [33]. Similarly, SII experiments typically implemented frequency binning through multiple narrow-band filters and standalone photon detectors, thereby limiting scalability [35, 36, 37, 11].
Other existing techniques, which could achieve simultaneously good spectral and temporal resolutions for the frequency multiplexing employ dispersive fibers to translate changes in time delay to the photon frequency [38], spectral phase shaping using spectral shear via electro-optic modulation [39], and an approach based on chirped fiber Bragg gratings [40]. These systems are interesting in their own right, but typically require complex infrastructure for pulsed operation and synchronization, and observation of the two-photon interference was not their primary motivation.
Maximizing interference contrast, or equivalently, visibility, necessitates simultaneously achieving exceptional temporal ($\sim$10 ps) and spectral ($\sim$10 pm) resolutions, thereby approaching the fundamental limits set by the Heisenberg Uncertainty Principle [41], also known as the Gabor or Fourier or Heisenberg-Gabor limits [42]. Current advances in multi-pixel single-photon detector arrays enable building of wideband spectrometers and also address challenges related to temporal resolution, which commonly limits the achievable contrast due to peak smearing in HBT measurements. These innovative data-driven detectors significantly enhance the practical applicability of quantum optical phenomena across various real-world scenarios, opening numerous opportunities for cutting-edge applications.
@@ 2.1 LinoSPAD2 detector
The key part of our spectrometer is the LinoSPAD2 camera. In the LinoSPAD2 sensor [45, 46], each pixel is a single-photon sensitive avalanche photodiode (SPAD) with a $26.2\times 26.2$ µm2 size and 25.1% fill factor. The sensor consists of a linear $512\times 1$ pixel array with a median dark count rate of $\sim 125$ cps/pixel, cross-talk of 0.2% [47, 48] and final photon detection efficiency of about 30% at 520 nm for the sensor equipped with microlenses [45, 49] while operating at room temperature. After calibrations, the temporal resolution was determined to be equal to 40 ps (rms) employing an SPDC source of photon pairs [50]. In this work, we employed 256 pixels (one half of the sensor). LinoSPAD2 has already demonstrated good performance in detecting the HBT photon bunching signatures in our prior experiments with neon and argon spectral lines [44, 34, 47].
@@ 2.2 Spectrometer
The light entering the spectrometer is polarized and filtered via a $(640\pm 5)$ nm bandpass filter before entering a fiber port. A 1-to-2 50:50 fiber-coupled beamsplitter splits the light into two beams, directing it to the two arms of the dual spectrometer. The outputs of the beamsplitter are connected to adjustable collimators that help with focusing and improving the spectral resolution. Each beam is then guided onto a ruled reflective diffraction grating with 1200 grooves/mm after reflection in a silver-coated mirror. The diffracted linear image of the spectrum within each arm of the dual spectrometer is focused through a 200 mm focal length lens onto the sensor. Every fiber used in this setup, including the fiber-coupled beamsplitter, is single-mode, and everything except the light sources is located inside a light-tight enclosure. To separate a possible cross-talk from the HBT effect, one of the two signal photons is delayed via an additional 1 m fiber in one of the dual spectrometer arms. This results in a shift of $\sim 5$ ns of the HBT peak relative to $\Delta t=0$ in the coincidence histograms [47, 48].
@@ 3.1 HBT peaks
Figure 4: a–d) Time difference distributions in two spectral bins in the dual spectrometer, four combinations in total, with examples of different contrast values for HBT peaks determined by the fit. e) Sum of all diagonal combinations where the two frequencies are expected to be equal. The peak contrast as determined by a fitting procedure is equal to $(2.0\pm 0.1)$% and sigma 70 ps; f) HBT peak after filtering with $(656.40\pm 0.05)$ nm narrow band filter.
To estimate the interference contrast in an unbiased manner, we developed a simple fitting procedure with a minimal number of free parameters and applied it uniformly to all combinations of pixel pairs. The HBT peak was fitted with a Gaussian function with its position predicted by the offset calibration [48] and constrained to $\pm 200$ ps of its nominal value. The Gauss sigma was constrained to the values from 60 to 80 ps, as its possible interval was determined in independent high-statistics measurements. The normalized background was set equal to one. The height of the peak — or contrast — was constrained to the values from -10 to 10%.
@@ 4 Discussion
Astrometric resolution: The presented results are an important step forward for the SII and quantum-assisted schemes of optical interferometers [52, 24] because the spectrometer can be directly employed for spectral binning in the corresponding experiments. In the case of the tested spectrometer, we can combine about 200 pairs of spectral bins where the contrast is nonzero, corresponding to three adjacent diagonal lines. The uncertainty of the opening angle between two stars $\Delta\theta$ can be expressed as in [14], showing dependence on the most important variables:
@@
It is proportional to the wavelength $\lambda$ and inversely proportional to the baseline between the interferometer stations $B$, allowing us to gain from longer baselines. The temporal and spectral resolutions of the spectrometer are accounted for by the visibility $V$, while other photon collection and detection parameters, such as the size of the mirrors and the detection efficiency, are accounted for in the detected rate of photons $\bar{n}$. With $T$ being the length of observation period we also note that the overall $T^{-3/2}$ dependence is more favorable than simply the expected $T^{-1/2}$ gain from photon pair statistics, reflecting the advantage of being able to use the measurement of a fringe oscillation rate due to the Earth rotation as proposed in [14].
The contrasts presented in Figure 5 have direct correspondence to $V$ and therefore in the presented study, we directly measured the visibility for each pair of spectral bins and can determine the projected uncertainty $\sigma\left[\Delta\theta\right]$ by combining measurements in multiple bins. Employing 70 spectral bins from the main diagonal allows to decrease the astrometric uncertainty by a factor of 8.3, while using extra bins from two adjacent diagonals, which are expected to have non-zero visibility, improved this value to 10.0. We consider this as an important step to demonstrate the advantages of frequency multiplexing and the viability of the approach. We also note that the number of spectral bins can be scaled up by a large factor employing larger and multiple sensors, and that it is not unusual to have tens of thousands of spectral channels on modern astronomical spectrographs [53].
@@
where $\hbar$ is Planck’s reduced constant, $\Delta E$, $\Delta f$, and $\Delta t$ are respectively the standard deviations for measurements of energy, frequency, and time. In our case of 640 nm wavelength and assuming 40 ps & 40 pm rms resolutions [44], $\Delta f\Delta t$ is a factor of 14.7 larger than the Heisenberg-Gabor limit, interpreting the spectral resolution uncertainty as frequency uncertainty. In general, this comparison is a useful benchmark for the indistinguishability of photons if they hit the same spectral bins in the two spectrometers. Experimentally, this means that the contrast of the corresponding HBT peak cannot be larger than $100\%/14.7=6.8\%$. The measured contrast of about 3-4% is indeed limited by the finite spectral and temporal resolutions of the spectrometer.
@@
Future advancements in fast sensor technologies are expected to significantly enhance temporal resolution, potentially achieving resolutions as fine as 5 ps rms for SPAD devices, as recently demonstrated in [58]. Similarly, superconducting nanowire single-photon detectors (SNSPDs) are rapidly advancing toward temporal resolutions in the single-digit picosecond regime. However, these detectors currently require cryogenic temperatures, and their scalability to multi-channel and two-dimensional detector arrays has not yet been fully established [59, 60, 61].
@@
5 Conclusion
We have demonstrated the first observation of the HBT effect for broadband light dispersed in a fast spectrometer across a wide, continuous spectral range of 10 nm, utilizing a large number, of the order of 100, of spectral channels. After precise calibration, our spectrometer achieved exceptional temporal and spectral resolutions of 40 ps and 40 pm, respectively. This enabled the observation of thermal photon bunching with an average contrast of approximately 3%, consistent with theoretical expectations. The achieved temporal and spectral filtering approaches the Heisenberg–Gabor limit, making our spectrometer highly applicable to diverse fields such as quantum communications, quantum metrology, spectroscopy, and quantum-assisted telescopy.
Analyzing the spectrometer’s performance specifically in the context of phase-sensitive intensity interferometry, we found that combining about 200 spectral bin combinations exhibiting non-zero contrast could enhance astrometric precision by a factor of 10. Furthermore, we have outlined the significant potential gains achievable through entanglement swapping in quantum networks, where frequency multiplexing could dramatically enhance the throughput and robustness of entangled photon distribution.
Future advancements in temporal and spectral resolution promise further significant improvements, potentially enabling near-ideal visibility of two-photon interference phenomena. Ultimately, our realization of massively frequency-multiplexed two-photon correlation measurements presents a transformative advancement, with the potential to accelerate the development of quantum sensing and quantum communication technologies by orders of magnitude.

### v2

@@ Abstract
Two-photon interference is a fundamental resource for quantum technologies and optical quantum computing, underpinning precision measurements, scalable entanglement distribution, and the operation of photonic circuits and quantum network protocols. Here, we report the first demonstration of massively parallel, wavelength-resolved photon bunching, revealing Hanbury Brown-Twiss correlations across 100 independent spectral channels. These observations are enabled by a fast, data-driven single-photon spectrometer that achieves 40 pm spectral and 40 ps temporal resolution over a 10 nm bandwidth, providing simultaneous access to spectro-temporal photon correlations without the need for narrowband filtering. This approach preserves photon flux while enabling high-dimensional quantum interference measurements across a broad spectrum. Our results establish frequency-multiplexed two-photon interference as a scalable and throughput-efficient platform for quantum-enhanced photonic technologies, offering a practical route toward room-temperature architectures that overcome loss limitations and advance the scalability for a variety of applications.
@@ 1 Introduction
Two-photon correlation analysis stands as a cornerstone of contemporary quantum optics, tracing its origins back to foundational studies by Glauber [23, 24]. This powerful method has found diverse applications across numerous fields, including super-resolution microscopy [17, 52], quantum communication protocols [57], two-photon resonance fluorescence [75, 68], stellar intensity interferometry (SII) [5, 19, 56, 31, 27], and its advanced phase-sensitive variants [25, 61, 2, 65]. Central to these applications are distinct correlation signatures between photons arising from fundamental quantum interference effects, predominantly characterized by the indistinguishability of photons as in the Hanbury Brown-Twiss (HBT) [4, 6] and Hong-Ou-Mandel (HOM) [30] phenomena.
Quantum interference is also a fundamental building block of optical quantum computing, underlying gate operations, photonic circuits, and quantum network protocols that connect distributed quantum processors [36, 60]. However, scalability in these systems is critically limited by optical loss, which constrains both communication distance and computational fidelity. A promising strategy to mitigate these limitations is multiplexing, where multiple parallel attempts — such as simultaneous Bell-state measurements — improve entanglement generation rates and overall network reliability. In this context, massively parallel, wavelength-multiplexed two-photon interference provides a practical route toward scalable, room-temperature quantum photonic architectures. By enabling many concurrent correlation measurements, this approach supports measurement-based quantum computing and quantum repeater networks, both of which rely on efficient entanglement swapping.
In this study, we directly measure and quantify the indistinguishability of the photons employing a fast spectrometer combined with the precise analysis of spectral and temporal data performed during post-processing, as schematically presented in Figure 1. For the first time, we unambiguously observe the HBT effect of photon bunching in broadband light simultaneously across 100 spectral bins in a continuous 10 nm range, establishing a new methodology for massively multiplexed correlation sensing over an extensive frequency range, with the potential of substantially improving the resolution, sensitivity, and scalability for quantum optical measurements.
@@
Another application for the spectrometer with excellent spectral and temporal resolutions is resonance fluorescence, where precise two-photon interference and correlation measurements can yield groundbreaking insights into quantum optical processes in atomic and molecular physics. Spectral correlation patterns observed in resonance fluorescence typically manifest as distinct photon bunching lines and anti-bunching circles, unveiling quantum behaviour at off-resonant frequencies [50, 75].
@@
Previous attempts to measure the frequency-multiplexed HBT effect were constrained by either limited temporal or spectral resolution or by a narrow frequency range with few spectral bins. Earlier investigations of spectrally resolved HOM interference employed slower cameras, utilizing nanosecond-scale resolution to identify temporal coincidences in photon pairs. Temporal coherence was usually maintained by precisely delaying one photon relative to the other using mechanical stages, effectively enabling sub-picosecond precision and thereby providing good visibility for HOM interference [73, 62, 71]. More recent efforts, such as those described in [18, 59, 21], achieved significantly improved timing resolution, 20–40 ps. However, these studies still faced limitations, with only a small number of spectral channels, eight channels in [18] and five narrow neon spectral lines in [21], or restricted spectral coverage, less than 0.3 nm, when investigating two-photon transitions in polariton condensates [59]. Similarly, SII experiments typically implemented frequency binning through multiple narrow-band filters and standalone photon detectors, thereby limiting scalability [69, 42, 67, 31].
Other existing techniques, which could achieve simultaneously good spectral and temporal resolutions for the frequency multiplexing, employ dispersive fibers to translate changes in time delay to the photon frequency [12], spectral phase shaping using spectral shear via electro-optic modulation [15], and an approach based on chirped fiber Bragg gratings [14]. These systems are interesting in their own right, but typically require complex infrastructure for pulsed operation and synchronization, and do not provide multi-photon detection capabilities as observation of the two-photon interference was not their primary motivation.
Maximizing interference contrast, or equivalently, visibility, necessitates simultaneously achieving exceptional temporal ($\sim$10 ps) and spectral ($\sim$10 pm) resolutions, thereby approaching the fundamental limits set by the Heisenberg Uncertainty Principle [28], also known as the Gabor or Fourier or Heisenberg-Gabor limits [22]. Current advances in multi-pixel single-photon detector arrays enable building of wideband spectrometers and also address challenges related to temporal resolution, which commonly limits the achievable contrast due to peak smearing in HBT measurements. These innovative data-driven detectors significantly enhance the practical applicability of quantum optical phenomena across various real-world scenarios by enabling simultaneous spectro-temporal two-photon correlation measurements without narrowband filtering. The presented approach provides a generally applicable experimental framework that is relevant not only for the mentioned quantum networks and stellar intensity interferometry, but also for a range of related applications, including thermal-light range sensing, pseudo-natural-light (anti)superbunching and ghost-imaging protocols, as well as studies of superthermal photon statistics in bimodal nanolasers [74, 47, 70, 64].
@@ 2.1 LinoSPAD2 detector
The key part of our spectrometer is the LinoSPAD2 camera. In the LinoSPAD2 sensor [8, 49], each pixel is a single-photon sensitive avalanche photodiode (SPAD) with a $26.2\times 26.2$ µm2 size and 25.1% fill factor. The sensor consists of a linear $512\times 1$ pixel array with a median dark count rate of $\sim 125$ cps/pixel, cross-talk of 0.2% [39, 40] and final photon detection efficiency of about 30% at 520 nm and 20% at 640 nm for the sensor equipped with microlenses [8, 7] while operating at room temperature. After calibrations, the temporal resolution was determined to be equal to 40 ps (rms) employing an SPDC source of photon pairs [41]. In this work, we employed 256 pixels (one half of the sensor). LinoSPAD2 has already demonstrated good performance in detecting the HBT photon bunching signatures in our prior experiments with neon and argon spectral lines [32, 21, 39].
@@ 2.2 Spectrometer
The light entering the spectrometer is polarized and filtered via a $(640\pm 5)$ nm bandpass filter before entering a fiber port. A 1-to-2 50:50 fiber-coupled beamsplitter splits the light into two beams, directing it to the two arms of the dual spectrometer. The outputs of the beamsplitter are connected to adjustable collimators that help with focusing and improving the spectral resolution. Each beam is then guided onto a ruled reflective diffraction grating with 1200 grooves/mm after reflection in a silver-coated mirror. The diffracted linear image of the spectrum within each arm of the dual spectrometer is focused through a 200 mm focal length lens onto the sensor.
Every fiber used in this setup, including the fiber-coupled beamsplitter, is single-mode, and everything except the light sources is located inside a light-tight enclosure. The total system efficiency (including the detector) after entering the fiber port was estimated to be about 10%.
To separate a possible cross-talk from the HBT effect, one of the two signal photons is delayed via an additional 1 m fiber in one of the dual spectrometer arms. This results in a shift of $\sim 5$ ns of the HBT peak relative to $\Delta t=0$ in the coincidence histograms [39, 40].
@@ 3.1 HBT peaks
Figure 4: a–d) Time difference distributions in two spectral bins in the dual spectrometer, four combinations in total (pixels 51 and 218, 90 and 179, 50 and 219, 46 and 223), with examples of different contrast values for HBT peaks determined by the fit. e) Sum of all diagonal combinations where the two frequencies are expected to be equal. The peak contrast as determined by a fitting procedure is equal to $(2.0\pm 0.1)$% and sigma 70 ps; f) HBT peak after filtering with $(656.40\pm 0.05)$ nm narrow band filter.
To estimate the interference contrast in an unbiased manner, we developed a simple fitting procedure with a minimal number of free parameters and applied it uniformly to all combinations of pixel pairs. The HBT peak was fitted with a Gaussian function with its position predicted by the offset calibration [40] and constrained to $\pm 200$ ps of its nominal value. The Gauss sigma ($\sigma$) was constrained to the values from 60 to 80 ps, as its possible interval was determined in independent high-statistics measurements. The normalized background was set equal to one. The height of the peak — or contrast ($C$) — was constrained to the values from -10 to 10%.
@@ 4 Discussion
Astrometric resolution: The presented results are an important step forward for the stellar intensity interferometry and quantum-assisted schemes of optical interferometers [13, 3] because the spectrometer can be directly employed for spectral binning in the corresponding experiments. In the case of the tested spectrometer, we can combine about 200 pairs of spectral bins where the contrast is non-zero, corresponding to three adjacent diagonal lines. In the considered two-photon interferometer the opening angle between two stars $\Delta\theta$ is extracted from population of HBT peaks measured for different combinations of detectors in two stations of the interferometer [61]. The associated uncertainty of the opening angle can be expressed as in [61], showing dependence on the most important variables:
@@
It is proportional to wavelength $\lambda$ and inversely proportional to baseline (i.e. distance) between interferometer stations $B$, allowing us to gain from longer baselines. The temporal and spectral resolutions of the spectrometer are accounted for by the visibility $V$, while other photon collection and detection parameters, such as the size of the mirrors and the detection efficiency, are accounted for in the detected rate of photons $\bar{n}$. With $T$ being the length of observation period we also note that the overall $T^{-3/2}$ dependence is more favorable than simply the expected $T^{-1/2}$ gain from photon pair statistics, reflecting the advantage of being able to use the measurement of a fringe oscillation rate due to the Earth rotation as proposed in [61] and elaborated in [13].
The contrasts presented in Figure 5 have the same definition as $V$. In the presented study, we directly measured the visibility for each pair of spectral bins and, therefore, can determine the projected uncertainty $\sigma\left[\Delta\theta\right]$ by combining visibility measurements in multiple bins, assuming that all other parameters in Eq. 4.1 are the same for different spectral bins. Employing 70 spectral bins from the main diagonal would allow us to decrease the astrometric uncertainty by a factor of 8.3, while using extra bins from two adjacent diagonals, which are expected to have non-zero visibility, improved this value to 10.0. We consider this as an important step to demonstrate the advantages of frequency multiplexing and the viability of the approach for this application. We also note that the number of spectral bins can be scaled up by a large factor employing larger and multiple sensors, and that it is not unusual to have tens of thousands of spectral channels on modern astronomical spectrographs [48].
@@
where $\hbar$ is Planck’s reduced constant, $\Delta E$, $\Delta f$, and $\Delta t$ are respectively the standard deviations for measurements of energy, frequency, and time.
In our case of 640 nm wavelength and assuming 40 ps & 40 pm rms resolutions [32], $\Delta f\Delta t$ is a factor of 14.7 larger than the Heisenberg-Gabor limit, interpreting the spectral resolution uncertainty as frequency uncertainty. In general, this comparison is a useful benchmark for the indistinguishability of photons if they hit the same spectral bins in the two spectrometers. Experimentally, this means that the contrast of the corresponding HBT peak cannot be larger than $100\%/14.7=6.8\%$. The measured contrast of about 3-4% is indeed limited by the finite spectral and temporal resolutions of the spectrometer.
@@
Current state of the art and future advancements: In Table 1 we compare our results with those of other published fast spectrometers, focusing on the operational wavelength $\lambda$, spectral and temporal resolutions $\Delta\lambda$ & $\Delta t$, multiplication factor with respect to the Heisenberg-Gabor limit (i.e. how far the product of resolutions is above the limit), number of spectral channels, and the possibility to measure multiple photons in parallel. All examples in the table are far inferior to our results in terms of the product of two resolutions and, in some cases, in the number of spectral channels.
Table 1: Comparison of fast single-photon spectrometers. Column "HGL" corresponds to the multiplication factor for the Heisenberg-Gabor limit, i.e. how far the product of resolutions is above the limit.
$\lambda$ (nm) $\Delta\lambda$ (nm) $\Delta t$ (ns) HGL # of chan. multi-photon ref.
640 0.04 0.04 14.7 256 yes this work
531 1.7 0.02 486 8 yes [18]
780 0.33 0.38 781 5 yes [67]
600 7.0 0.03 2420 200 no [11]
500 2.0 0.11 3300 32 yes [33]
800 0.5 7.30 22300 230 yes [20]
800 0.7 6.00 24700 256 yes [72]
An important case for comparison is presented in [11], which employed a different approach. It used a single channel of the long superconducting nanowire single-photon detector (SNSPD) with a diffraction grating to disperse light along the nanowire, determining the wavelength from the time difference of signal arrival at its two ends and leveraging the detector’s excellent timing resolution. However, the spectral resolution achieved was much poorer than in our work, with long and irregular tails in the resolution function.
Future advancements in fast sensor technologies are expected to significantly enhance temporal resolution, potentially achieving resolutions as fine as 5 ps rms for SPAD devices, as recently demonstrated in [26]. Hybrid solutions with a SPAD sensor attached to a time-stamping chip are also being investigated as a route to large two-dimensional fast imagers [44, 29, 45, 9]. Similarly, SNSPDs are rapidly advancing toward temporal resolutions in the single-digit picosecond regime. However, these detectors currently require cryogenic temperatures, and their scalability to multi-channel and two-dimensional detector arrays has not yet been fully established [10, 37, 54].
@@
Data availability
The raw measured data set is over 2 TB in size, and the preprocessed data that is used for histogram building is also approximately 2 TB in size, making it difficult to share online. Nonetheless, the data is archived and can be provided upon reasonable request.